OpenStudy (anonymous):

The length of a rectangle is 2 inches less than three times its width, and its perimeter is 36 inches. Find the width of the rectangle.

OpenStudy (bahrom7893):

woops sorry wrong it's actually: L = 3W - 2 P = 2 ( L+W )

OpenStudy (bahrom7893):

Now plug in for L P = 2 ( 3W - 2 + W ) = 36

OpenStudy (bahrom7893):

2 ( 4W - 2 ) = 36 8W - 4 = 36 8W = 36 + 4 8W = 40 W = 5

OpenStudy (bahrom7893):

L = 3W - 2 L = 3*5 - 2 L = 15-2 L = 13